I am struggling to figure out which formula to use to solve this problem. Any help is appreciated. Thanks in advance.
Jim is going to invest to help with a down payment on a home. How much would he have to invest to have $49,600 after 9 years, assuming an interest rate of 1.69% compounded annually?
Do not round any intermediate computations, and round your final answer to the nearest dollar.

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

formula is f=p*(1+r)^n
f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.

he needs to have 49600 after 9 years, so f = 49600.
interest rate is 1.69% compounded annually, so r = .0169 per year.
n = 9 years
formula becomes:
49600 = p*(1+.0169)^9
divide both sides of the equation by ((1+.0169)^9) to get:
49600/((1+.0169)^9) = 42655.73705.
replace p in the original equation to get:
f = 42655.73705*(1+.0169)^9 = 49600.
this confirms the solution is correct.

your answer is that he will have to invest 42655.74 today so that he will have 49600 in 9 years.


Answer by ikleyn(52933)

.

To see many similar solved problems, see the lessons
    - Compound interest percentage problems
    - Problems on discretely compound accounts
in this site, and learn the subject from there.


After reading these lessons, you will tackle such problems on your own without asking for help from outside.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Logarithms".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.


Happy learning (!)